# Jean DELHAYE

### v2

@@ 2.2. The cutoff phenomenon
One may consider the universal enveloping C*-algebra $C(U_{N}^{+})$ associated with $\mathcal{O}(U_{N}^{+})$ (see, for instance, [7, Sec II.8.3]). By definition, any state on $\mathcal{O}(U_{N}^{+})$ uniquely extends to a state on $C(U_{N}^{+})$, thus yielding an element of the Fourier-Stieltjes algebra, which corresponds to the topological dual $C(U_{N}^{+})^{*}$ of $C(U_{N}^{+})$. By the Riesz representation theorem, this dual space can be viewed as a noncommutative analogue of the measure algebra, equipped with the total variation norm. We denote the norm on this dual space as $\|\cdot\|_{FS}$, referring to it as the Fourier-Stieltjes norm.
Furthermore, the topological double dual $C(U_{N}^{+})^{**}$ of $C(U_{N}^{+})$ is recognized as the universal enveloping von Neumann algebra of $C(U_{N}^{+})$, serving as a noncommutative and universal analogue of classical measure spaces. Using the theory of Haagerup’s noncommutative $\mathrm{L}^{p}$-spaces, we can readily adapt the argument from [21, Lem 2.6] to show that:
$\frac{1}{2}\|\varphi-\psi\|_{FS}=\sup_{p\in\mathcal{P}}|\varphi(p)-\psi(p)|,$
where $\mathcal{P}$ denotes the set of orthogonal projections in $C(U_{N}^{+})^{**}$, which can be thought of as the noncommutative counterparts to indicator functions of Borel subsets. The proof details are omitted here, as they are not required for our current discussion. This generalized total variation distance will be employed in our analysis of cutoff phenomena, and we will denote this distance as
$d_{\mathrm{TV}}(\varphi,\psi)=\frac{1}{2}\|\varphi-\psi\|_{FS}.$
In the classical setting, a particularly significant case arises when both $\mu$ and $\nu$ are absolutely continuous with respect to the Haar measure. We can consider a similar scenario in the quantum context. Define an inner product on $\mathcal{O}(U_{N}^{+})$ by $\langle x,y\rangle=h(xy^{*})$. Completing this inner product space yields the Hilbert space $\mathrm{L}^{2}(U_{N}^{+})$, and $\mathcal{O}(U_{N}^{+})$ embeds into $\mathrm{B}(\mathrm{L}^{2}(U_{N}^{+}))$ via left multiplication (see [29, Cor 1.7.5] and the comments thereafter). The weak closure of this image is denoted by $\mathrm{L}^{\infty}(U_{N}^{+})$ and forms a von Neumann algebra.
If $\varphi:\mathcal{O}(U_{N}^{+})\to\mathbb{C}$ is a linear map that extends to a normal bounded map on $\mathrm{L}^{\infty}(U_{N}^{+})$, then $\varphi$ is an element of the Fourier algebra, which is the Banach space predual $\mathrm{L}^{\infty}(U_{N}^{+})_{*}$ of $\mathrm{L}^{\infty}(U_{N}^{+})$. In this case, $\|\varphi\|_{FS}=\|\varphi\|_{\mathrm{L}^{\infty}(U_{N}^{+})_{*}}$ (see [11, Prop 3.14]), which further implies, by [21, Lem 2.6], that:
$d_{\mathrm{TV}}(\varphi,\psi)=\sup_{p\in\widetilde{\mathcal{P}}}|\varphi(p)-\psi(p)|,$
where $\widetilde{\mathcal{P}}$ is the set of orthogonal projections in $\mathrm{L}^{\infty}(U_{N}^{+})$. It is important to note that for this formula to be applicable, the states $\varphi$ and $\psi$ must extend to the von Neumann algebra $\mathrm{L}^{\infty}(U_{N}^{+})$.
@@
Let $(\mathbb{G}_{N},\varphi^{(N)}_{t\geq 0})_{N\in\mathbb{N}}$ be a sequence of compact quantum groups each equipped with a Lévy process. We say that they exhibit a cutoff phenomenon at time $(t_{N})_{N\in\mathbb{N}}$ if for any $\epsilon>0$ we have
@@
where $h_{N}$ denotes the Haar state of $\mathbb{G}_{N}$. More precisely, given a continuous function $f$ decreasing from $1$ to 00 such that
@@
Note that a limit profile is unique only up to affine transformation. Specifically, suppose a given limit profile $f$ is attained along the sequences $(t_{N},s_{N})_{N}$. Then, for any $a>0$ and $b\in\mathbb{R}$, the limit profile $c\mapsto f(ac+b)$ is achieved along the sequences $(t_{N}+bs_{N},as_{N})_{N}$. Now, consider a single process that exhibits two distinct limit profiles, $f$ and $f^{\prime}$, along the sequences $(t_{N},s_{N})_{N}$ and $(t^{\prime}_{N},s^{\prime}_{N})_{N}$, respectively. Up to an affine transformation, we may assume that these limit profiles coincide at two points, say 00 and $1$, and that $f$ is strictly decreasing in a neighborhood of both 00 and $1$ (as the function decreases from 1 to 0). Then, if $d_{N}(t)$ denotes the distance between the process and the Haar state at time $t$, we have:
@@
Given that the $d_{N}$’s are decreasing (see [23, Lem 2.6]), it follows that $t^{\prime}_{N}=t_{N}+o(s_{N})$ and $s^{\prime}_{N}=s_{N}+o(s_{N})$. To see this, suppose, for instance, that the first equality does not hold. This would imply that, possibly after passing to a subsequence, we have
@@
In this paper, we explore the limit profile of the Brownian motion on $U_{N}^{+}$ which is shared by the Brownian motion on $O_{N}^{+}$, with the primary challenge being the identification of suitable sequences $(t_{N})_{N\in\mathbb{N}}$ and $(s_{N})_{N\in\mathbb{N}}$. We denote by $\nu_{\mathrm{SC}}$ the semi-circular distribution, which is defined as follows:
@@
and its $\mathrm{L}^{2}$-space has the Chebyshev polynomials as an orthonormal basis, as introduced just before Theorem 2.4. For any real number $c$, let $\eta_{c}$ represent the distribution defined by
$\mathrm{d}\eta_{c}=(1-e^{2c})\mathbf{1}_{c<0}\delta_{e^{c}+e^{-c}}+\frac{e^{c}\sqrt{4-x^{2}}\mathbf{1}_{|x|<2}}{2\pi(e^{c}+e^{-c}-x)}\mathrm{d}x.$
In essence, $\eta_{c}$ is the shifted free Meixner law $\mathrm{Meix}(-e^{-c},0)*\delta_{e^{-c}}$. For further details on free Meixner laws, we refer the reader to [8, Sec 2.2]. It is important to note that the semi-circular distribution can be considered a free Meixner law with parameters $(0,0)$, emphasizing that $\eta_{c}$ is effectively a deformation of $\nu_{\mathrm{SC}}$. Moreover, it has been established in [23, Lem 3.12] that $\eta_{c}$ is the unique measure satisfying
$\eta_{c}(P_{n})=e^{-cn},\quad n\in\mathbb{N}.$
The limit profile of the Brownian motion on $U_{N}^{+}$ is determined by the function mapping each $c\in\mathbb{R}$ to the total variation distance between $\eta_{c}$ and $\nu_{\mathrm{SC}}$.
@@ 3.1. Defining the quantum unitary Brownian motion
It is not entirely clear which generating functional on $U_{N}^{+}$ could serve as the analogue of the Brownian motion. Fortunately, a well-defined concept of Brownian motion on $O_{N}^{+}$ has been established. F. Cipriani, U. Franz, and A. Kula provided a classification of central generating functionals on $O_{N}^{+}$ in [12, Thm 10.2] using an analogue of the Lévy-Khinchine formula. Specifically, they determined that all such functionals have the following form:
@@
for some $b\geq 0$ and positive Borel measure $\nu$ on $[-N,N)$.
@@
Comparing this formula with the one established by M. Liao for classical compact Lie groups in [25], we see that the process corresponding to $\nu=0$ plays a role analogous to the one associated to the Laplace-Beltrami operator. As a consequence, this functional is called the Brownian motion on $O_{N}^{+}$.
@@
We lack a similar decomposition for central generating functionals on $U_{N}^{+}$, making it challenging to define Brownian motion on this quantum group. Fortunately, the Brownian motion on $O_{N}^{+}$ has other properties that we can leverage to define a Brownian motion on $U_{N}^{+}$.
@@
Firstly, any centralized Gaussian generating functional on $O_{N}^{+}$ qualifies as a Brownian motion (see the Appendix for a proof). Secondly, as proved in [10, Prop 3.9], the Laplace-Beltrami operator on $O_{N}$ acts as a Brownian motion on $O_{N}^{+}$ when composed with the quotient map $\mathcal{O}(O_{N}^{+})\to\mathcal{O}(O_{N})$.
[…]
$U_{w}:=U_{w_{1}}\otimes...\otimes U_{w_{n}}\quad\text{and}\quad\chi_{w}:=\chi_{U_{w}},$
where $U_{\lozenge}=U$ and $U_{\blacklozenge}=\overline{U}$. We also denote by $p(w)$, $q(w)$, and $\ell(w)$ the number of occurrences of $\lozenge$, $\blacklozenge$, and the total number of letters in the word $w$, respectively. We claim that functionals $L:\mathcal{O}(U_{N}^{+})_{0}\to\mathbb{C}$, satisfying for any word $w$:
(3.2) $L_{w}:=L(\chi_{w})=-\ell N^{\ell-1}\alpha-\Big{(}{(p-q)^{2}-\ell}\Big{)}N^{\ell-2}\beta,$
where $p:=p(w)$, $q:=q(w)$, and $\ell:=\ell(w)$, for some $\alpha,\beta\geq 0$ with $\alpha\geq\beta/N$, adhere to the expected behavior and properties of Brownian motions, thereby justifying their classification as such.
As expected, centralized real-valued Gaussian generating functionals on $U_{N}^{+}$ are of this form (see the Appendix for a proof). Let us prove that the Laplace-Beltrami operator on $U_{N}$ is of this form when considered on $U_{N}^{+}$.
@@
The Lie algebra $\mathfrak{u}_{N}$ of the unitary group $U_{N}$ consists of $N\times N$ skew-Hermitian matrices. The family
$X_{ij}=\left\{\begin{array}[]{cc}{E_{ij}-E_{ji}/2\sqrt{N}}&\text{if }i<j,\\ {\mathrm{i}E_{ii}/\sqrt{2(N-1)}}&\text{if }i=j,\\ {\mathrm{i}(E_{ij}+E_{ji})/2\sqrt{N}}&\text{if }i>j,\end{array}\right.\quad 1\leq i,j\leq N,$
where $E_{ij}$ is the matrix with a 1 in the $(i,j)$-th entry and zeros elsewhere, forms an orthonormal basis for the negative Killing form of $\mathfrak{u}_{N}$, defined by
$B(X,Y)=-2N\operatorname{Tr}(XY)+2\operatorname{Tr}(X)\operatorname{Tr}(Y).$
@@
We will use $X_{ij}$ to refer to both the matrix and the associated differential operator. The Laplace-Beltrami operator on $U_{N}$ is defined as $L=(\sum_{i,j}X_{ij}^{2})\big{|}_{g=e}$. We will always consider $L$ through the quotient map $\mathcal{O}(U_{N}^{+})\to\mathcal{O}(U_{N})$ and conditional expectation $\mathbb{E}:\mathcal{O}(U_{N}^{+})_{0}\to\mathcal{O}(U_{N}^{+})$.
@@
[…]
$L(\chi_{w})=-\ell(w)N^{\ell(w)-1}\alpha-\Big{(}(p(w)-q(w))^{2}-\ell(w)\Big{)}N^{\ell(\beta)-2}\beta,$
with $\alpha=(N^{2}-N+1)/2(N-1)$ and $\beta=N/2(N-1)$.
@@
Let us compute.
$L_{\circ}:=L(\chi_{U})$ $=\sum_{i,j}X_{ij}^{2}\operatorname{Tr}|_{e}=\sum_{i,j}\frac{\partial^{2}}{\partial s\partial t}\left.\operatorname{Tr}\left(\exp(sX_{ij})\exp(tX_{ij})\right)\right|_{t,s=0}$
$=\sum_{i,j}\operatorname{Tr}(X_{ij}^{2})=-\frac{N^{2}-N+1}{2(N-1)}=-\alpha;$
$L_{\circ\circ}:=L(\chi_{U\otimes U})$ $=\sum_{i,j}X_{ij}^{2}(\operatorname{Tr}\cdot\operatorname{Tr})|_{e}$
$=\sum_{i,j}X_{ij}\left.\left(X_{ij}(\operatorname{Tr})\operatorname{Tr}+\operatorname{Tr}X_{ij}(\operatorname{Tr})\right)\right|_{e}$
$=\sum_{i,j}\left.\left(X_{ij}^{2}(\operatorname{Tr})\operatorname{Tr}+2X_{ij}(\operatorname{Tr})X_{ij}(\operatorname{Tr})+\operatorname{Tr}X_{ij}^{2}(\operatorname{Tr})\right)\right|_{e}$
$=2NL(\chi_{U})+2\sum_{i}\operatorname{Tr}(X_{ii})^{2}=2N\alpha-\frac{N}{N-1}$
$=-2N\alpha-2\beta;$
$L_{\circ\bullet}:=L(\chi_{U\otimes\overline{U}})$ $=\sum_{i,j}X_{ij}^{2}(\operatorname{Tr}\cdot\overline{\operatorname{Tr}})|_{e}=\sum_{i,j}X_{ij}\left.\left(X_{ij}(\operatorname{Tr})\overline{\operatorname{Tr}}+\operatorname{Tr}X_{ij}(\overline{\operatorname{Tr}})\right)\right|_{e}$
$=\sum_{i,j}\left.\left(X_{ij}^{2}(\operatorname{Tr})\overline{\operatorname{Tr}}+2X_{ij}(\operatorname{Tr})X_{ij}(\overline{\operatorname{Tr}})+\operatorname{Tr}X_{ij}^{2}(\overline{\operatorname{Tr}})\right)\right|_{e}$
$=N(L(\chi_{U})-L(\chi_{\overline{U}}))+2\sum_{i,j}\operatorname{Tr}(X_{ij})\overline{\operatorname{Tr}}(X_{ij})=-2N\alpha+\frac{N}{N-1}$
$=-2N\alpha+2\beta.$
where $\overline{\operatorname{Tr}}$ denotes the function
$\overline{\operatorname{Tr}}:U_{N}\longrightarrow\mathbb{C},g\longmapsto\overline{\operatorname{Tr}(g)}.$
This shows the result for words of length $\leq 2$. Assume now that $w=w_{1}...w_{\ell}$ is a word of length $\ell\geq 3$ and set $p:=p(w)$ and $q:=q(w)$. Let us fix some notations. We will write $W$ to denote the map $\operatorname{Tr}_{w_{1}}\cdot...\cdot\operatorname{Tr}_{w_{n}}$ where $\operatorname{Tr}_{\lozenge}=\operatorname{Tr}$ and $\operatorname{Tr}_{\blacklozenge}=\overline{\operatorname{Tr}}$. We will also write $W_{n}$ (resp. $W_{mn}$) to denote the product map of the $\operatorname{Tr}_{w_{j}}$’s omitting $\operatorname{Tr}_{w_{n}}$ (resp. omitting $\operatorname{Tr}_{w_{m}}$ and $\operatorname{Tr}_{w_{n}}$). Let us finally compute.
$L(\chi_{w})$ $=\sum_{i,j}X_{ij}^{2}W|_{e}=\sum_{i,j}X_{ij}\left.\left(\sum_{n}X_{ij}(\operatorname{Tr}_{w_{n}})W_{n}\right)\right|_{e}$
$=\sum_{i,j}\left.\left(\sum_{n}X_{ij}^{2}(\operatorname{Tr}_{w_{n}})W_{n}+\sum_{m\neq n}X_{ij}(\operatorname{Tr}_{w_{m}})X_{ij}(\operatorname{Tr}_{w_{n}})W_{mn}\right)\right|_{e}$
$=-\sum_{n}W_{n}(e)\psi_{w_{n}}+\sum_{i,j}\sum_{m\neq n}W_{mn}(e)\operatorname{Tr}_{w_{m}}(X_{ij})\operatorname{Tr}_{w_{n}}(X_{ij})$
$=N^{\ell-1}(pL_{\lozenge}+qL_{\blacklozenge})-N^{\ell-2}\sum_{m\neq n}\frac{N(L_{w_{m}}+L_{w_{n}})-L_{w_{m}w_{n}}}{2}$
$=N^{\ell-2}\left(\frac{p(p-1)}{2}L_{\lozenge\lozenge}+\frac{q(q-1)}{2}L_{\blacklozenge\blacklozenge}+pqL_{\lozenge\blacklozenge}\right)-N^{\ell-1}(\ell-2)\left(pL_{\lozenge}+qL_{\blacklozenge}\right)$
$=-{\ell}N^{\ell-1}\alpha-\Big{(}{(p-q)^{2}-\ell}\Big{)}N^{\ell-2}\beta.$

### v3

@@ 2.2. The cutoff phenomenon
Let us introduce a distance between states on $U_{N}^{+}$ that serves as a quantum analogue of the total variation distance. To motivate this definition, let us briefly recall the classical setting. Given a compact group $G$, its associated algebra of regular functions, denoted $\mathcal{O}(G)$, consists of polynomial functions in the matrix coefficients of finite-dimensional representations of $G$. The total variation distance between two probability measures $\mu$ and $\nu$ on $G$ is given by
$d_{\mathrm{TV}}(\mu,\nu)=\sup_{A\subset G}\left|\mu(A)-\nu(A)\right|=\sup_{A\subset G}\left|\int_{G}\mathbf{1}_{A}\mathrm{d}\mu-\int_{G}\mathbf{1}_{A}\mathrm{d}\nu\right|,$
where the supremum is taken over all Borel subsets of $G$. The issue is that, apart from the trivial ones, indicator functions $\mathbf{1}_{A}$ do not belong to $\mathcal{O}(G)$, but rather to a larger function space – they exactly correspond to orthogonal projections in $\mathrm{L}^{\infty}(G)$. We therefore seek an appropriate extension of $\mathcal{O}(U_{N}^{+})$ within a noncommutative $\mathrm{L}^{\infty}$-space. Following the classical approach, we equip $\mathcal{O}(U_{N}^{+})$ with an inner product induced by the Haar state, given by
$\langle x,y\rangle:=h(xy^{*}),$
which defines the Hilbert space $\mathrm{L}^{2}(U_{N}^{+})$. The left regular representation of $\mathcal{O}(U_{N}^{+})$ on $\mathrm{L}^{2}(U_{N}^{+})$, given by left multiplication (see [NT13, Cor 1.7.5] and the comments thereafter), gives raise to a von Neumann algebra, denoted by $\mathrm{L}^{\infty}(U_{N}^{+})$, as the weak closure of its image. This provides a natural setting in which to define a quantum analogue of the total variation distance. For two states $\varphi$ and $\psi$ on $\mathrm{L}^{\infty}(U_{N}^{+})$, we define their total variation distance as
$d_{\mathrm{TV}}(\varphi,\psi)=\sup_{p}\left|\varphi(p)-\psi(p)\right|,$
where the supremum is taken over all orthogonal projections in $\mathrm{L}^{\infty}(U_{N}^{+})$. The issue in the quantum setting is that a state on $\mathcal{O}(U_{N}^{+})$ may not necessarily extend to a normal state on $\mathrm{L}^{\infty}(U_{N}^{+})$ (namely when absolute continuity w.r.t. the Haar state is lost). A resolution, given in [FTW21], is to consider the universal enveloping $C^{*}$-algebra $C(U_{N}^{+})$ (see [Bla06, Sec II.8.3]), where any state on $\mathcal{O}(U_{N}^{+})$ uniquely extends to a state on $C(U_{N}^{+})$, yielding an element of the Fourier-Stieltjes algebra $C(U_{N}^{+})^{*}$, a noncommutative analogue of the measure algebra. Denoting by $\|\cdot\|_{\mathrm{FS}}$ the Fourier-Stieltjes norm, one can show using arguments from [Fre19, Lem 2.6] that for any two states $\varphi,\psi$ on $U_{N}^{+}$,
$\frac{1}{2}\|\varphi-\psi\|_{\mathrm{FS}}=\sup_{p}|\varphi(p)-\psi(p)|,$
where the supremum is taken over all orthogonal projections in $C(U_{N}^{+})^{**}$, the topological double dual of $C(U_{N}^{+})$, serving as the analogue of the measure space. If $\varphi$ and $\psi$ extend to states on $\mathrm{L}^{\infty}(U_{N}^{+})$, this supremum coincides with that over projections in $\mathrm{L}^{\infty}(U_{N}^{+})$ (see [BR17, Prop 3.14] and [Fre19, Lem 2.6]), so that we may extend the definition:
$d_{\mathrm{TV}}(\varphi,\psi):=\frac{1}{2}\|\varphi-\psi\|_{\mathrm{FS}}.$
@@
Let $(\mathbb{G}_{N},(\varphi^{(N)}_{t})_{t\geq 0})_{N\in\mathbb{N}}$ be a sequence of compact quantum groups each equipped with a Lévy process. We say that they exhibit a cutoff phenomenon at time $(t_{N})_{N\in\mathbb{N}}$ if for any $\epsilon>0$ we have
@@
where $h_{N}$ denotes the Haar state of $\mathbb{G}_{N}$. More precisely, given a continuous function $f$ decreasing from $1$ to 0 such that
@@
Note that a limit profile is unique only up to affine transformation. Specifically, suppose a given limit profile $f$ is attained along the sequences $(t_{N},s_{N})_{N}$. Then, for any $a>0$ and $b\in\mathbb{R}$, the limit profile $c\mapsto f(ac+b)$ is achieved along the sequences $(t_{N}+bs_{N},as_{N})_{N}$. Now, consider a single process that exhibits two distinct limit profiles, $f$ and $f^{\prime}$, along the sequences $(t_{N},s_{N})_{N}$ and $(t^{\prime}_{N},s^{\prime}_{N})_{N}$, respectively. Up to an affine transformation, we may assume that these limit profiles coincide at two points, say 0 and $1$, and that $f$ is strictly decreasing in a neighborhood of both 0 and $1$ (as the function decreases from 1 to 0). Then, if $d_{N}(t)$ denotes the distance between the process and the Haar state at time $t$, we have:
@@
Given that the $d_{N}$’s are decreasing (see [FTW21, Lem 2.6]), it follows that $t^{\prime}_{N}=t_{N}+o(s_{N})$ and $s^{\prime}_{N}=s_{N}+o(s_{N})$. To see this, suppose, for instance, that the first equality does not hold. This would imply that, possibly after passing to a subsequence, we have
@@
In this paper, we explore the limit profile of the Brownian motion on $U_{N}^{+}$. We denote by $\nu_{\mathrm{SC}}$ the semicircle distribution, which is defined as follows:
@@
and its $\mathrm{L}^{2}$-space has the (rescaled) Chebyshev polynomials as an orthonormal basis, as introduced just before Theorem 2.4. For any real number $c$ and $r\in[0,\infty]$, let $\eta_{c}$ represent the distribution defined by
$\eta_{c}^{r}(A)=\mathbb{E}\left[\left(1-e^{2c}R_{r}^{-2}\right)_{+}\mathbf{1}_{A}(e^{-c}R_{r}+e^{c}R_{r}^{-1})\right]+\int_{A}\mathbb{E}\left[\frac{1}{e^{2c}R_{r}^{2}-xe^{-c}R_{r}+1}\right]\mathrm{d}\nu_{\mathrm{SC}}(x),$
where $R_{r}=\cos T_{r}$ with $T_{r}$ being a random variable of Gaussian distribution $\mathcal{N}(0,2r)$ with the conventions that $T_{0}=0$ and $T_{\infty}\sim\mathrm{Unif}([0,2\pi])$. Note that $\eta_{c}^{0}$ is a shifted free Meixner law:
$\eta_{c}^{0}=(1-e^{2c})_{+}\delta_{e^{c}+e^{-c}}+\frac{\mathrm{d}\nu_{\mathrm{SC}}(x)}{e^{2c}-xe^{-c}+1}=\operatorname{Meix}^{+}(-e^{-c},0)*\delta_{e^{-c}}.$
For further details on free Meixner laws, we refer the reader to [BB06, Sec 2.2]. In that regard, $\eta_{c}^{r}$ appears as a mixing measure of the family
$\frac{\mathrm{d}\nu_{\mathrm{SC}}(x)}{q^{2}-xq+1}+(1-q^{2})_{+}\delta_{q+q^{-1}}=\operatorname{Meix}^{+}(-q,0)*\delta_{q},\quad q\in\mathbb{R}^{*}.$
It is important to note that the semicircle distribution can be considered a free Meixner law with parameters $(0,0)$, emphasizing that $\eta_{c}^{r}$ is effectively a deformation of $\nu_{\mathrm{SC}}$. Moreover, it has been established in [FTW21, Lem 3.12] that $\eta_{c}^{0}$ is the unique measure satisfying
$\eta_{c}^{0}(P_{n})=e^{-cn},\quad n\in\mathbb{N}.$
Remark 2.9.
Note that one easily checks that the orthogonal family of monic polynomials for $\mathrm{L}^{2}(\eta_{c}^{0})$ is given by
$\widehat{P}_{n}=P_{n}-e^{-c}P_{n-1},\quad n\in\mathbb{N},$
with the convention that $P_{-1}=0$. By inspecting this family and using [SY01, Thm 2.1], one obtains the result from [FTW21, Lem 3.12] without using free cumulants.
The limit profile of the Brownian motion on $U_{N}^{+}$ is determined by the function mapping each $c\in\mathbb{R}$ to the total variation distance between $\eta_{c}^{r}$ and $\nu_{\mathrm{SC}}$ [^1] Brownian motions on $U_{N}^{+}$ will emerge as a two parameter family, the constant $r$ essentially translates the quotient of those parameters..
@@ 3.1. Defining the quantum unitary Brownian motion
It is not entirely clear which generating functional on $U_{N}^{+}$ could serve as the analogue of the Brownian motion. Fortunately, a well-defined concept of Brownian motion on $O_{N}^{+}$ has been established. F. Cipriani, U. Franz, and A. Kula provided a classification of central generating functionals on $O_{N}^{+}$ in [CFK14, Thm 10.2] using an analogue of the Lévy-Khinchine formula. Specifically, they determined that all such functionals have the following form:
@@
for some $b\geq 0$ and positive Borel measure $\nu$ on [N,N)[-N,N).
@@
Comparing this formula with the one established by M. Liao for classical compact Lie groups in [Lia04], we see that the process corresponding to $\nu=0$ plays a role analogous to the one associated to the Laplace-Beltrami operator. As a consequence, this functional is called the Brownian motion on $O_{N}^{+}$.
@@
We lack a similar decomposition for central generating functionals on $U_{N}^{+}$, making it challenging to define the Brownian motion on this quantum group. Fortunately, the Brownian motion on $O_{N}^{+}$ has other properties that we can leverage to define a Brownian motion on $U_{N}^{+}$:
  1. (i)
    Centralized Gaussian generating functionals on $O_{N}^{+}$ are exactly the Brownian motions on $O_{N}^{+}$ (see Appendix A for a proof and a clear definition as to what is a Gaussian generating functional in the quantum setting).
  2. (ii)
@@
    Secondly, as proved in [BGJ20, Prop 3.9], the Laplace-Beltrami operator on $O_{N}$ acts as a Brownian motion on $O_{N}^{+}$ when composed with the quotient map $\mathcal{O}(O_{N}^{+})\to\mathcal{O}(O_{N})$ and centralized.
We will start from that last observation. Let us set some notations, we write $U=(u_{ij})_{1\leq i,j\leq N}$ and $\overline{U}=(u_{ij}^{*})_{1\leq i,j\leq N}$. For a word $w=w_{1}\cdots w_{n}$ over the set $\{\lozenge,\blacklozenge\}$, we define:
$U_{w}:=U_{w_{1}}\otimes\cdots\otimes U_{w_{n}}\quad\text{and}\quad\chi_{w}:=\chi_{U_{w}},$
where $U_{\lozenge}=U$ and $U_{\blacklozenge}=\overline{U}$. We also write:
  • $\ell(w)=n$ the length of the word;
  • $p(w)/q(w)$ the number of occurrences of $\lozenge/\blacklozenge$ in the word.
@@
The Lie algebra $\mathfrak{u}_{N}=\mathfrak{su}_{N}\oplus\mathfrak{u}_{1}$ of the unitary group $U_{N}$ consists of $N\times N$ skew-Hermitian matrices. Consider an orthonormal basis $(X_{i})_{1\leq i<N^{2}}$ of $\mathfrak{su}_{N}=\mathfrak{u}_{N}\cap\ker\operatorname{Tr}$ for the inner product
$\langle X,Y\rangle=-\operatorname{Tr}(XY),$
which we complete with
$X_{N^{2}}:=\frac{\mathrm{i}I_{N}}{\sqrt{N}}\in\mathrm{i}I_{N}\mathbb{R}\simeq\mathfrak{u}_{1}$
to obtain a full basis of the Lie algebra. We will allow for the abuse of notation and for any $X\in\mathfrak{u}_{N}$, denote by $X$ both the matrix and the differential operator acting on regular (even smooth) functions $f\in\mathcal{O}(U_{N})$, defined by
$Xf(x):=\left.\frac{\mathrm{d}}{\mathrm{d}t}f\left(x\exp(tX)\right)\right|_{t=0}.$
The Laplace-Beltrami operator is defined as:
$L(f):=\sum_{i=1}^{N^{2}}X_{i}^{2}f(e),\quad f\in\mathcal{O}(U_{N}^{+}).$
@@
It is a generating functional and we will always consider it through the quotient map $\mathcal{O}(U_{N}^{+})\to\mathcal{O}(U_{N})$ and conditional expectation $\mathbb{E}:\mathcal{O}(U_{N}^{+})_{0}\to\mathcal{O}(U_{N}^{+})$.
Remark 3.1.
Given a semisimple compact Lie group, one usually defines the Laplace–Beltrami operator via an orthonormal basis w.r.t. its negative Killing form. The group $U_{N}$ is not semisimple, and its negative Killing form, defined by
$B(X,Y)=-2N\,\operatorname{Tr}(XY)+2\,\operatorname{Tr}(X)\operatorname{Tr}(Y),\quad X,Y\in\mathfrak{u}_{N},$
is degenerate, as $\mathfrak{u}_{1}\not\subseteq\ker\operatorname{Tr}$.
@@
If $L$ denotes the Laplace-Beltrami operator on $U_{N}$, $\pi$ the quotient map $\mathcal{O}(U_{N}^{+})\to\mathcal{O}(U_{N})$ and $\mathbb{E}$ the conditional expectation $\mathcal{O}(U_{N}^{+})\to\mathcal{O}(U_{N}^{+})_{0}$, then the generating functional $\widehat{L}:=L\circ\pi\circ\mathbb{E}$ is defined by
$\widehat{L}(\chi_{w})=-\Big(\ell(w)\alpha+\big(p(w)-q(w)\big)^{2}\beta\Big)N^{\ell(w)-1}$
with $\alpha=N^{2}-1$ and
@@
Let us denote by $\operatorname{Tr}_{\lozenge}$ and $\operatorname{Tr}_{\blacklozenge}$ the linear forms $g\mapsto\operatorname{Tr}(g)$ and $g\mapsto\overline{\operatorname{Tr}(g)}$, respectively. Let us first compute:
$L_{\lozenge}$ $:=L(\operatorname{Tr}_{\lozenge})=\sum_{i=1}^{N^{2}}\operatorname{Tr}(X_{i}^{2})=-N^{2},$
$L_{\lozenge\lozenge}$ $:=L(\operatorname{Tr}_{\lozenge}\cdot\operatorname{Tr}_{\lozenge})=\sum_{i=1}^{N^{2}}2\left(\operatorname{Tr}(e)\operatorname{Tr}(X_{i}^{2})+\operatorname{Tr}(X_{i})^{2}\right)=-2N^{3}-2N,$
$L_{\lozenge\blacklozenge}$ $:=L(\operatorname{Tr}_{\lozenge}\cdot\operatorname{Tr}_{\blacklozenge})=\sum_{i=1}^{N^{2}}\left(\operatorname{Tr}(e)\overline{\operatorname{Tr}(X_{i}^{2})}+\overline{\operatorname{Tr}(e)}\operatorname{Tr}(X_{i}^{2})+\left|\operatorname{Tr}(X_{i})\right|^{2}\right)$
$=-2N^{3}+2N.$
This shows that $L$ is of the desired form for
$\alpha=N^{2}-1\quad\text{and}\quad\beta=1,$
at least for elements $\chi_{w}$, where $w$ is a word of length $\leq 2$. Let $w=w_{1}\cdots w_{n}$ be a word of length $n\geq 3$, with $p$ occurrences of $\lozenge$ and q=npq=n-p occurrences of $\blacklozenge$. Let us denote by $W$ the product map of the $\operatorname{Tr}_{w_{j}}$’s, by $W_{k}$ the same product omitting $\operatorname{Tr}_{w_{k}}$, and by $W_{k\ell}$ the product omitting both $\operatorname{Tr}_{w_{k}}$ and $\operatorname{Tr}_{w_{\ell}}$. We may compute:
$L(\chi_{w})$ $=\sum_{i}X_{i}^{2}W\big|_{e}=\sum_{i}X_{i}\left.\left(\sum_{k}X_{i}(\operatorname{Tr}_{w_{k}})W_{k}\right)\right|_{e}$
$=\sum_{i}\left.\left(\sum_{k}X_{i}^{2}(\operatorname{Tr}_{w_{k}})W_{k}+\sum_{k\neq\ell}X_{i}(\operatorname{Tr}_{w_{\ell}})X_{i}(\operatorname{Tr}_{w_{k}})W_{k\ell}\right)\right|_{e}$
$=\sum_{k}W_{k}(e)L_{w_{k}}+\sum_{i}\sum_{k\neq\ell}W_{k\ell}(e)\operatorname{Tr}_{w_{k}}(X_{i})\operatorname{Tr}_{w_{\ell}}(X_{i})$
$=-nN^{n+1}-N^{n-2}\Bigg(pq\left(L_{\lozenge\blacklozenge}-N(L_{\lozenge}+L_{\blacklozenge})\right)+\frac{p(p-1)+q(q-1)}{2}\left(L_{\lozenge\lozenge}-2NL_{\lozenge}\right)\Bigg)$
$=-nN^{n+1}-N^{n-1}\left(p(p-1)+q(q-1)-2pq\right)$
$=-\left(n(N^{2}-1)+(p-q)^{2}\right)N^{n-1}.$